A 100 cm cube solution of sodium carbonate is prepared by dissolving 8.653g of the sail in water. The density of solution is 1.0816 pee millilitres. What are the molarity and molality of the solution.

Grade:12

3 Answers

Arun
6 years ago
Dear Akash

Density = mass / volume ;
1.0816 = mass / 100 ;  mass=108.16
mass of solvent = 108.16 - 8.653 = 99.507
Molar mass of sodium carbonate = 23 + 12 + 3*16 = 83
no. of moles of sodium carbonate = (given mass) / molar mass=8.653 / 83 = 0.104
molality = ( no. of moles ) / ( mass of solvent ) = ( 0.104 / 99.507 ) * 1000 = 1.0451
molarity = ( no. of moles ) / ( volume of solution ) = ( 0.104 / 100 ) * 1000=1.04
 
Regards
Arun (askIITians forum expert)
Mahesh
3 years ago
Bhai aap ko sodium carbonate ka formula bhi nahii ataa. Aap Akash ka naam kharab kar rahe hooo.
 
Molar mass Na2 CO3 = 23x2 + 12+ 16x3
                                       = 46 + 12+48
                                        = 106g/mol
Aditya singh
2 years ago
D=m/v 
1.0816=m/100, m=1.0816×100=108.16g
Mass of solvent =108.16-8.653= 99.507
No.of moles of solute = 8.653/106=0.0816
M=no.of moles of solute/volume of solution in l
M =0.0816mol×1000/100cm3= 0.816M
Molality(m)=no.of moles of solute /mass of solvent inkg 
m= 0.0816 mol×1000/99.507g= 0.820m